The table shows the number of people living in a house and the weight of trash (in pounds) at the curb just before the trash pickup.

People(x) Trash(y) Zx Zy
2 18 -0.4392 -0.6838
3 33 0.1464 -0.2519
5 93 1.3175 1.4756
1 23 -1.0247 -0.5399
Answer #1
X Y X^2 Y^2 XY
2 18 4 324 36
3 33 9 1089 99
5 93 25 8649 465
1 23 1 529 23
Total 11 167 39 10591 623

SS_x = \sum X^2 - \frac{(\sum X)^2}{n} = 39-\frac{11^2}{4} = 8.75

SS_y = \sum Y^2 - \frac{(\sum Y)^2}{n} = 10591-\frac{167^2}{4} = 3618.75

SS_xy = \sum XY - \frac{\sum X \sum Y}{n} = 623-\frac{167*11}{4} = 163.75

r = \frac{SS_xy }{\sqrt{SS_ySS_x}}

r = \frac{163.75 }{\sqrt{3618.75*8.75}}

r = 0.9202